# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ 3.2 Norming sets for polynomials
In view of Lemma 1, the unisolvency of the polynomial interpolation problem is equivalent to the invertibility of the corresponding Vandermonde matrix. Building on it, in this section, we recall the notion of norming sets, which establishes qualitative relations among the properties of the polynomials, the minimal singular value of the Vandermonde matrices and spatial dispersion of the set $A$.
Let $\Omega$ be a compact subset of $\mathbb{R}^{d}$. Then equipping the class of $l$th order polynomials $\mathscr{P}_{l}(\Omega)$ with the supremum norm
$\|P\|=\sup_{x\in\Omega}|P(x)|,\;\forall P\in\mathscr{P}_{l}(\Omega),$
forms a normed vector space. We say that a finite set $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\Omega$ is a norming set for $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}>0$ if
@@
The notion of norming sets was originally introduced by [24] for general normed vector spaces and their duals. In this paper we focus on polynomials restricted to a compact set and point evaluation functionals on this compact set. Note that for a norming set $A$, the norm $\|\cdot\|$ and the weaker norm $\|\cdot\|_{A}$ defined as
@@
are equivalent over $\mathscr{P}_{l}(\Omega)$. Heuristically, this means that the behaviour of a polynomial $P\in\mathscr{P}_{l}(\Omega)$ on a norming set $A\subset\Omega$ characterizes its behavior on the whole compact set $\Omega$. The existence of such finite set follows from the finite, $l+d\choose l$} dimensionality of the vector space $\mathscr{P}_{l}(\Omega)$. The following lemma connects norming sets with Vendermonde matrices.
Lemma 2.
Let $\Omega=[0,1]^{d}$. For $l\in\mathbb{N}$, and set $A=\{w_{1},w_{2},\cdots,w_{m}\}$ with cardinality $m={l+d\choose l}$, the following two statements are equivalent:
  1. (1)
    The set $A$ is a norming set on $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}$, for some $c_{N}>0$;
  2. (2)
    The minimal singular value of the Vandermonde matrix $V_{A}$ is bounded from below by $c_{S}$, for some $c_{S}>0$.
Moreover, there exists a universal constant $C_{0}\geq 1$, not depending on $A$, such that
$C_{0}^{-1}c_{N}\leq c_{S}^{-1}\leq C_{0}c_{N}.$
While the equivalence between norming sets and unisolvent sets is well known in the literature, the above lemma characterizes the qualitative relationship between the constants $c_{N}$ and $c_{S}$. This explicit dependence plays a crucial role in computing the small deviation bounds for Vecchia GPs in Section 5. We note, that an important difference between Lemma 1 and Lemma 2 is the support of the polynomials. In view of the unboundedness of polynomials on the whole domain $\mathbb{R}^{d}$, finite norming sets do not exist for $\Omega=\mathbb{R}^{d}$.
The norming constant $c_{N}$ has an explicit expression and is “scale-invariant”, see [8].
Lemma 3.
Let $\Omega$ be a compact subset of $\mathbb{R}^{d}$ and let $A\subset\Omega$ with cardinality $m={l+d\choose l}$ be a norming set of $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}$. Then we have
$c_{N}=\sup_{x\in\Omega}\|V_{A}^{-1}v_{x}\|_{1}.$ (17)
Moreover, for all $\tau>0$, the set $\tau A$ is a norming set of $\mathscr{P}_{l}(\tau\Omega)$ with the same norming constant $c_{N}$.
However, the above explicit formula (17) for the norming constant $c_{N}$, can still be challenging to compute, i.e. one has to invert the corresponding Vandermonde matrix and take supremum over the whole compact set $\Omega$. For the unidimensional case, i.e. $\Omega=[0,1]$, we provide an alternative formulation. Let $A=\{w_{1},w_{2},\cdots,w_{l+1}\}\subset[0,1]$ consists of $l+1$ distinct points. Then $c_{N}$ can be expressed with the help of Lagrangian polynomials, i.e.
$c_{N}=\sup_{x\in[0,1]}\sum_{j=1}^{l+1}\bigg{|}\frac{\prod_{i\neq j}(x-w_{i})}{\prod_{i\neq j}(w_{j}-w_{i})}\bigg{|}.$ (18)
In higher dimensions, however, without further, technical assumptions on the set $A$, we can not simplify equation (17).
[…]
First, note that as the order $l$ of the polynomials and hence the cardinality of the corresponding norming sets increases, the norming constants also increase. This is intuitively clear, since the size of the polynomial space increases with the order, hence it is becoming more difficult to characterize them with a finite set of functionals. Second, one can observe that for randomly chosen points $w_{j}$, the norming constant is substantially higher than for the equidistant design.
Then we consider the $d=2$ case, see Figure 2. Similarly to the one dimensional case one can notice the increase of the norming constants in the order of the polynomials. Furthermore, the norming constants are smaller for the so called corner sets than for i.i.d. samples generated from uniform distribution. The corner set is a carefully designed unisolvent set capturing the local polynomial structure, see Section 4.2 for its definition and further discussion. Moreover, for $d\geq 2$, unlike in the one dimensional case, not all set with cardinality $l+d\choose l$} is a norming set for $\mathscr{P}_{l}(\Omega)$. We provide these examples at the bottom row of Figure 2.
Figure 1: Norming sets for a unidimensional space for polynomials of order $1,2,3$ and $4$, respectively. Top row: Equally spaced points on a $[0,1]$; bottom row: i.i.d. samples from the uniform distribution on $[0,1]$.
Figure 2: Norming sets in $0,1]^{2}$. The cardinality of the sets in the $l$th ($1\leq l\leq 4$) column is $l+2\choose l$}, which is the dimension of the linear space $\mathscr{P}_{l}(\Omega)$. Top row: “corner sets”; middle row: i.i.d. samples from the uniform distribution; bottom row: non-norming sets for $\mathscr{P}_{l}(\Omega)$.
4 Layered Norming DAGs
After discussing some background material on polynomial interpolation in the previous section, we focus here on the DAG structure used in the Vecchia approximation.
@@ 4.1 The Layered and Norming conditions
We start by introducing the layered and norming conditions, which in turn define the Layered Norming DAGs, used in our proposal. For simplicity, we consider the domain $\mathcal{X}=[0,1]^{d}$.
We start by describing the Layered condition.
@@
  • $\forall X_{i}\in\mathcal{N}_{j}$, $\mathrm{pa}(X_{i})\subset\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$;
  • There exist constants $\gamma>1$, $c_{d}>0$, such that $\forall X_{i_{1}},X_{i_{2}}\in\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$, $\|X_{i_{1}}-X_{i_{2}}\|_{\infty}\geq c_{d}\gamma^{-j}$.
In the rest of the paper, we assume that the indeces of the vertices in $\mathcal{X}_{n}$ are ordered with respect to their corresponding layers. More precisely, denoting by $\eta(i)$ the layer of $X_{i}$, i.e. $X_{i}\in\mathcal{N}_{\eta(i)}$, then for all $X_{i},X_{i^{\prime}}\in\mathcal{X}_{n}$ with $i<i^{\prime}$, we have $\eta(i)\leq\eta(i^{\prime})$.
The Layered condition, as the name suggests, partitions the training set $\mathcal{X}_{n}$ into $J+1$ layers $\mathcal{N}_{0},\mathcal{N}_{1},\ldots,\mathcal{N}_{J}$, such that for any $X_{i}\in\mathcal{N}_{j}$, the associated parents can only belong to layers preceding $\mathcal{N}_{j}$. Moreover, the minimum distance between the elements is governed by their respective layers. The overall design of the layered structure facilitate a multi-resolution framework, where the first few layers tend to be relatively spread out in the domain while the latter layers gradually fill in the gaps of the domain. The layered condition is very mild and can be satisfied for arbitrary training set $\mathcal{X}_{n}$. Next we introduce the Norming condition for DAGs.
@@
Consider a DAG satisfying Condition 1 with parameter $\gamma>1$. Let us take $m={l+d\choose l}$ and assume that there exists $i_{0}\in\mathbb{N}$, such that for all $i\geq i_{0}$, $|\mathrm{pa}(X_{i})|=m$. Moreover, there exist constants $c_{L},c_{N}>0$, such that for all $X_{i}\in\mathcal{N}_{j}$, with $0\leq j\leq J$ and $i\geq i_{0}$, there exists a $d$-dimensional cube $\mathcal{C}\supset\mathrm{pa}(X_{i})\cup\{X_{i}\}$ with side length no greater than $c_{L}\gamma^{-j}$, such that $\mathrm{pa}(X_{i})$ is a norming set on $\mathscr{P}_{l}(\mathcal{C})$ with norming constant $c_{N}$.
The above Norming condition, built upon the Layered condition, requires that the parent set of each $X_{i}$ (from a threshold $i_{0}$ onwards) is a norming set on a neighborhood of $X_{i}$. Moreover, it assumes the existence of a universal norming constant $c_{N}$ for these norming sets.
Unlike Condition 1 that can be satisfied for all $\mathcal{X}_{n}$ with a properly built DAG, the Norming condition, especially the requirement of a universal norming constant $c_{N}$, is not always straightforward. In the univariate case it does hold automatically, since any set consisting of $m$ distinct elements is a norming set for polynomial space of order m1m-1 as computed in equation (18). However, it is more complicated in the multivariate case, where not every set of size $m={l+d\choose l}$ forms a norming set for $\mathscr{P}_{l}([0,1]^{d})$, let alone with a universal norming constant. Nevertheless, even in the higher dimensional cases there exist specific sets, e.g. grid data or Fekete points [8], with universally bounded norming constants.
@@ 4.2 Building Layered Norming DAGs
In this section we provide a constructive algorithm for finding DAG structures that satisfies Conditions 1 and 2. As discussed in the previous subsection, the latter condition can be challenging in a high-dimensional space. For simplicity, we restrict our attention to the grid data, while deferring the discussion on general data set $\mathcal{X}_{n}$ to Section 8.1 and Section A.5 of the Supplementary Material.
@@
We first consider the case $\tilde{n}=2^{r}+1,r\in\mathbb{N}$. The construction of the Layered Norming DAGs consists of two steps, i.e. constructing layers and assigning parent sets. The layers $\mathcal{N}_{j},0\leq j\leq r$ are defined as follows. The base (root) layer $\mathcal{N}_{0}$ is defined as
@@
In other words, the layer $\mathcal{N}_{0}$ consists the $2^{d}$ vertices of the $d$-dimensional unite cube. For $1\leq j\leq r$, the layer $\mathcal{N}_{j}$ is defined as
@@
The layers $\mathcal{N}_{j}$ defined above satisfy Condition 1 with parameters $\gamma=2$ and $c_{d}=1$. It remains to specify the parent sets $\mathrm{pa}(X_{i})$ for all $X_{i}$ such that Condition 2 is also satisfied. Let $j_{0}$ be the smallest positive integer such that $2^{j_{0}-1}+1\geq l+1$. In other words, $j_{0}$ is the first layer such that there are sufficient number of elements in the preceding layers $\mathcal{N}_{0},\mathcal{N}_{1},\cdots,\mathcal{N}_{j_{0}-1}$ to build norming sets. The vertices in $\mathcal{N}_{0}$ have no parents. For all $X_{i}\in\mathcal{N}_{j},0<j<j_{0}$, the parent set is defined as $\mathrm{pa}(X_{i})=\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$. For vertices in $\mathcal{N}_{j},j\geq j_{0}$ we employ the corner sets discussed in Section 3.2 of [39], see below.
For all $X_{i}\in\mathcal{N}_{j}$, the corresponding parent set is a subset of
@@
For each coordinate $1\leq h\leq d$, we order the coordinates $\{k/2^{j-1}:0\leq k\leq 2^{j-1}\}$ according to their distance from $X_{i}[h]$. In the case of ties, elements tying can be ordered arbitrarily. Let us denote this ordered set as
@@
The corner parent set of $X_{i}$ is defined as
@@
The name corner comes from the formula $\sum_{h=1}^{d}[i_{h}-1]\leq l$, meaning the indices of coordinate values, in ascending order with respect to the distance to each $X_{i}[h]$, is at the corner of the $d$-dimensional grid. The shapes and norming constants of some corner sets are given in Figure 2.
Note that the proposed parent sets in (21) have cardinality $l+d\choose l$}. Furthermore, the corresponding polynomial interpolation is unisolvent and in view of the multivariate divided-difference formula, see Theorem 5 of [39], has an explicit form. Therefore, in view of Lemma 2, each of these parent sets is a norming set. Unfortunately, we can not write the norming constant as a simple and explicit function of $d$ and $l$. However, there are only a finite number of possible shapes for the corner sets, each possessing a finite norming constant. Taking their maximum results in an absolute norming constant over all possible corner sets, providing Condition 2.
Finally, for $\tilde{n}\neq 2^{r^{\prime}}+1$, $\forall r^{\prime}\in\mathbb{N}$, let $r$ be the largest integer such that $\tilde{n}>2^{r}+1$. Let the index set $\mathscr{I}\subset\{0,1,\cdots,\tilde{n}-1\}$ have cardinality $2^{r}+1$ such that $\{0,1,\cdots,\tilde{n}-1\}\backslash\mathscr{I}$ does not contain consecutive integers. Then we consider the subset of $\mathcal{X}_{n}$ defined by
@@
is on a $d$-dimensional grid with cardinality $(2^{r}+1)^{d}$ that has a one-to-one correspondance with $\bar{X}_{n}$ induced by the coordinate-wise mapping $\mathcal{M}$. Thus, one can build a Layered Norming DAG on $\hat{\mathcal{X}}_{n}$ with the previous discussed method, which will subsequently induce a Layered Norming DAG on $\bar{X}_{n}$. Once this is done, we let the layer $\mathcal{N}_{r+1}=\mathcal{X}_{n}\backslash\bar{\mathcal{X}}_{n}$. Then, for all $X_{i}\in\mathcal{N}_{r+1}$, we define the parent set to be a corner set as in equation (21). We formalize the above method in Algorithm 1, deferred to Section A.4 of the Supplementary Material. We also visualize our DAG on a $2$-dimensional grid, with the layer structures in Figure 3 and the parent sets in Figure 4.
@@ 5 Probabilistic properties
In this section, we study the probabilistic properties of the proposed Vecchia approximation of Matérn processes given in Section 1.3. The standard techniques for analyzing Matérn GPs heavily rely on the explicit form of the covariance kernel and its Fourier transform. Unfortunately, in case of Vecchia GPs these techniques can not be applied. First, the covariance matrices of their finite dimensional marginals can only be obtained by inverting the precision matrices, rendering the direct analysis substantially more difficult. Moreover, the covariance kernel of the Vecchia approximated Matérn process is not stationary, hence the techniques relying on the spectral density are not adequate anymore. This necessitates the development of new mathematical tools built on the finite dimensional conditional distributions of the GP.
The section is organized as follows. In Section 5.1 we collect some results regarding the derivatives of Matérn covariance functions. Section 5.2 studies the conditional distribution of $Z^{\tau,s}_{X}$ given the same process at $\mathrm{pa}(X)$. Section 5.3 utilizes the results of Section 5.2 to derive small deviation bounds for the Vecchia GP $\hat{Z}^{\tau,s}_{X}$ associated to the Matérn mother GP. Finally, Section 5.4 studies the Reproducing Kernel Hilbert Space (RKHS) associated with Vecchia GPs and their approximation properties. These results are used to derive lower bound for the de-centered small ball probabilities of Vecchia GPs. Throughout the section, whenever a DAG is involved, we assume that Conditions 1 and 2 are satisfied with some positive constants $c_{d},c_{L}$ and $c_{N}$ not depending on the sample size $n$.
The proofs for the lemmas and theorems in Section 5.1 and 5.2 are provided in Section B, while the proofs of the lemmas and theorems in Section 5.3 and 5.4 are deferred to Section C of the Supplementary Material.

### v4

@@ 3.1 Norming sets for polynomials
A norming set is a key concept that connects the properties of polynomials with the spatial dispersion of certain finite point sets. Let $\Omega$ be a compact subset of $\mathbb{R}^{d}$. Equipping the class of $l$th-order polynomials $\mathscr{P}_{l}(\Omega)$ with the supremum norm
$\|P\|=\sup_{x\in\Omega}|P(x)|,\;\forall P\in\mathscr{P}_{l}(\Omega)$
forms a normed vector space. A finite set $A=\{w_{1},w_{2},\cdots,w_{m}\}\subset\Omega$ is said to be a norming set for $\mathscr{P}_{l}(\Omega)$ with norming constant $c_{N}>0$ if
@@
The notion of norming sets was originally introduced by [25] for general normed vector spaces and their duals. In this paper, we focus on polynomials restricted to a compact set and point evaluation functionals on this set. Note that for a norming set $A$, the norm $\|\cdot\|$ and the weaker seminorm $\|\cdot\|_{A}$
@@
are equivalent on $\mathscr{P}_{l}(\Omega)$. Heuristically, this means that the behavior of a polynomial $P\in\mathscr{P}_{l}(\Omega)$ on a norming set $A\subset\Omega$ characterizes its behavior on the entire compact domain $\Omega$. The existence of such finite set follows from the finite, $l+d\choose l$} dimensionality of the vector space $\mathscr{P}_{l}(\Omega)$.
@@ 3.2 The Layered and Norming conditions
We begin by introducing the Layered and Norming conditions, which together define the layered norming DAGs, used in our proposal. For simplicity, we consider the domain $\mathcal{X}=[0,1]^{d}$.
We first describe the Layered condition.
@@
  • $\mathrm{pa}(X_{i})\subset\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$, $\forall X_{i}\in\mathcal{N}_{j}$;
  • $\exists$ $\gamma>1$, $c_{d}>0$, such that $\|X_{i_{1}}-X_{i_{2}}\|_{\infty}\geq c_{d}\gamma^{-j}$, $\forall X_{i_{1}},X_{i_{2}}\in\cup_{j^{\prime}=0}^{j}\mathcal{N}_{j^{\prime}}$.
Throughout the remainder of the paper, we assume that the indices of the vertices in $\mathcal{X}_{n}$ are ordered according to their corresponding layers. More precisely, denoting by $\eta(i)$ the layer of $X_{i}$, i.e. $X_{i}\in\mathcal{N}_{\eta(i)}$, we assume that for all $X_{i},X_{i^{\prime}}\in\mathcal{X}_{n}$ with $i<i^{\prime}$, we have $\eta(i)\leq\eta(i^{\prime})$.
The Layered condition, as the name suggests, partitions the training set $\mathcal{X}_{n}$ into $J+1$ layers $\mathcal{N}_{0},\mathcal{N}_{1},\ldots,\mathcal{N}_{J}$, such that for any $X_{i}\in\mathcal{N}_{j}$, its parents can only belong to layers preceding $\mathcal{N}_{j}$. Moreover, the minimum distance between elements is governed by their respective layers. This layered structure facilitates a multiresolution framework in which the first few layers tend to be relatively spread out over the domain, while the latter layers gradually fill in the gaps. The layered condition is very mild and can be satisfied for an arbitrary training set $\mathcal{X}_{n}$.
Next we introduce the Norming condition for DAGs.
@@
Consider a DAG satisfying Condition 1 with parameter $\gamma>1$. Let $m={l+d\choose l}$ and assume that there exists $i_{0}\in\mathbb{N}$, such that for all $i\geq i_{0}$, $|\mathrm{pa}(X_{i})|=m$. Moreover, there exist constants $c_{L},c_{N}>0$, such that for all $X_{i}\in\mathcal{N}_{j}$, with $0\leq j\leq J$ and $i\geq i_{0}$, there exists a $d$-dimensional cube $\mathcal{C}\supset\mathrm{pa}(X_{i})\cup\{X_{i}\}$ with side length no greater than $c_{L}\gamma^{-j}$, such that $\mathrm{pa}(X_{i})$ is a norming set on $\mathscr{P}_{l}(\mathcal{C})$ with norming constant $c_{N}$.
The above Norming condition, built upon the Layered condition, requires that the parent set of each $X_{i}$ (for $i\geq i_{0}$) is a norming set on a neighborhood of $X_{i}$. Moreover, it assumes the existence of a universal norming constant $c_{N}$ for these norming sets.
Unlike Condition 1, which can be satisfied for any $\mathcal{X}_{n}$ with a properly constructed DAG, the Norming condition, especially the requirement of a universal norming constant $c_{N}$, is not always straightforward to verify. In the univariate case it holds automatically, since any set of $m$ distinct points is a norming set for the polynomial space of order m1m-1, as computed in equation (B.9). However, the situation is more complicated in the multivariate case, where not every set of size $m={l+d\choose l}$ forms a norming set for $\mathscr{P}_{l}([0,1]^{d})$, let alone with a universal norming constant. Nevertheless, even in the higher dimensional cases there exist specific sets, such as grid design or Fekete points [8], that do possess uniformly bounded norming constants.
@@ 3.3 Building layered norming DAGs
In this section, we provide a constructive algorithm for finding DAG structures that satisfies Conditions 1 and 2. As discussed in the previous subsection, the latter condition can be challenging to meet in high-dimensional spaces. For simplicity, we restrict our attention to grid data, and defer the discussion of general data set $\mathcal{X}_{n}$ to Section 6.1 and Section C.2 of the Supplementary Material.
@@
We first consider the case $\tilde{n}=2^{r}+1,r\in\mathbb{N}$. The construction of the layered norming DAGs consists of two steps: constructing the layers and assigning the parent sets. The layers $\mathcal{N}_{j},0\leq j\leq r$ are defined as follows. The base (root) layer $\mathcal{N}_{0}$ is
@@
In other words, the layer $\mathcal{N}_{0}$ consists of the $2^{d}$ vertices of the $d$-dimensional unit cube. For $1\leq j\leq r$, the layer $\mathcal{N}_{j}$ is defined as
@@
The layers $\mathcal{N}_{j}$ defined above satisfy Condition 1 with parameters $\gamma=2$ and $c_{d}=1$. It remains to specify the parent sets $\mathrm{pa}(X_{i})$ for all $X_{i}$ such that Condition 2 is also satisfied. Let $j_{0}$ be the smallest positive integer such that $2^{j_{0}-1}+1\geq l+1$. In other words, $j_{0}$ is the first layer for which the preceding layers $\mathcal{N}_{0},\mathcal{N}_{1},\cdots,\mathcal{N}_{j_{0}-1}$ contain a sufficient number of points to form norming sets. The vertices in $\mathcal{N}_{0}$ have no parents. For all $X_{i}\in\mathcal{N}_{j},0<j<j_{0}$, the parent set is defined as $\mathrm{pa}(X_{i})=\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$. For vertices in $\mathcal{N}_{j},j\geq j_{0}$ we employ the corner sets discussed in Section 3.2 of [40], as described below.
For each $X_{i}\in\mathcal{N}_{j}$, the corresponding parent set is a subset of
@@
For each coordinate $1\leq h\leq d$, we order the values $\{k/2^{j-1}:0\leq k\leq 2^{j-1}\}$ by their distance from $X_{i}[h]$. In the case of ties, the tied elements may be ordered arbitrarily. Let us denote the resulting ordered set by
@@
The corner parent set of $X_{i}$ is then defined as
@@
The term corner refers to the condition $\sum_{h=1}^{d}[i_{h}-1]\leq l$, which ensures that the selected coordinate indices, ordered by increasing distance from each $X_{i}[h]$, lie near the “corner” of the $d$-dimensional index grid. Examples of the shapes and norming constants of several corner sets are given in Figure B.2 of the supplement.
Note that the proposed parent sets in (16) have cardinality $l+d\choose l$}. Moreover, the corresponding polynomial interpolation problem is unisolvent and by the multivariate divided-difference formula (see Theorem 5 of [40]), admits an explicit representation. Consequently, in view of Lemma B.2, each of such parent set is a norming set. Although the norming constant cannot be written as a simple explicit function of $d$ and $l$, there are only finitely many possible shapes for the corner sets, and each possess a finite norming constant. Taking their maximum yields an absolute norming constant that applies uniformly to all possible corner sets, thereby establishing Condition 2.
Finally, for $\tilde{n}\neq 2^{r^{\prime}}+1$, $\forall r^{\prime}\in\mathbb{N}$, let $r$ be the largest integer such that $\tilde{n}>2^{r}+1$. Let $\mathscr{I}\subset\{0,1,\cdots,\tilde{n}-1\}$ be an index set of cardinality $2^{r}+1$ such that $\{0,1,\cdots,\tilde{n}-1\}\backslash\mathscr{I}$ does not contain consecutive integers. Then we consider the subset of $\mathcal{X}_{n}$ defined by
@@
lies on a $d$-dimensional grid of cardinality $(2^{r}+1)^{d}$ and has a one-to-one correspondence with $\bar{X}_{n}$ induced by the coordinate-wise mapping $\mathcal{M}$. Thus, one can build a layered norming DAG on $\hat{\mathcal{X}}_{n}$ using the method described previously, which subsequently induces a layered norming DAG on $\bar{X}_{n}$. Once this is done, we define the layer $\mathcal{N}_{r+1}=\mathcal{X}_{n}\backslash\bar{\mathcal{X}}_{n}$. For all $X_{i}\in\mathcal{N}_{r+1}$ the parent set is defined as a corner set as in equation (16). We formalize this procedure in Algorithm 1, which is deferred to Section C.1 of the Supplementary Material. The DAG on a $2$-dimensional grid is visualized with the layer structures in Figure 1 and the parent sets in Figure 2.
@@ 4 Probabilistic properties
In this section, we study the probabilistic properties of the proposed Vecchia approximation of Matérn processes introduces in Section 1.3. Standard techniques for analyzing Matérn GPs rely heavily on the explicit form of the covariance kernel and its Fourier transform. Unfortunately, these techniques are not applicable to Vecchia GPs. First, the covariance matrices of their finite dimensional marginals can only be obtained by inverting the precision matrices, rendering the direct analysis substantially more difficult. Moreover, the covariance kernel of the Vecchia approximated Matérn process is not stationary, hence techniques based on spectral density are no longer adequate. This motivates the development of new mathematical tools that exploit the finite-dimensional conditional distributions of the GP.
The section is organized as follows. In Section 4.1, we collect several results regarding the derivatives of Matérn covariance functions. Section 4.2 studies the conditional distribution of $Z^{\tau,s}_{X}$ given the process values at $\mathrm{pa}(X)$. Section 4.3 utilizes the results from Section 4.2 to derive small deviation bounds for the Vecchia GP $\hat{Z}^{\tau,s}_{X}$ associated with the Matérn mother GP. Finally, Section 4.4 investigates the Reproducing Kernel Hilbert Space (RKHS) associated with Vecchia GPs and their approximation properties, which are then used to derive lower bounds for the decentered small ball probabilities of Vecchia GPs. Throughout the section, whenever a DAG is involved, we assume that Conditions 1 and 2 are satisfied with some positive constants $c_{d},c_{L}$ and $c_{N}$ that do not depend on the sample size $n$.